Statistics 2
Folien in diesem Satz (10)
-
- Factorials
- n objects in a row can be arranged in n! ways.
- n! = n(n – 1)(n – 2)(n – 3) × .... × 4 × 3 × 2 ×1
- Note that 0! is defined to be 1. This fits in with formulae for combinations.
- Combinations
- The number of ways we can choose r objects from a total of n objects, where the order does not
- matter, is called the number of combinations of r objects from n and is written as
- nCr = n(n-1)(n-2)...(n-r+1)/r! or nCr = n!/(n-r)!r!
- We can think of this as n choose r.
- Example: Find the number of hands of 4 cards which can be dealt from a pack of 10.
- Solution: In a hand of cards the order does not matter, so this is just the number of combinations of 4
- from 10, or 10 choose 4
- 10C4 = 10x9x8x7/4! = 210
-
- To write down the expansion of (p + q)
- 6
- We write down the terms in a logical order then use the numbers in the ‘6th’ row of the triangle.
- Binomial coefficients and combinations
- The binomial coefficients ቀ
- ݊
- ݎቁ are equal to the number of combinations nCr
- ⇒ = nCr = n!/(n-r)!r!
-
- Conditions for a Binomial Distribution
- - A single trial has exactly two possible outcomes – success and failure.
- - This trial is repeated a fixed number, n, times.
- - The n trials are independent.
- T
- - The probability of success, p, remains the same for each trial.
- The probability of success in a single trial is usually taken as p and the probability of failure as q.
- Note that p + q = 1.
-
Cumulative binomial probability tables
- Example: For X ~ B(30, 0⋅35), find the probability that 7
- Solution: A moment’s thought shows that we need
- P(X = 8, 9, 10, 11 or 12)
- = P(X ≤ 12) – P(X ≤ 7) = 0⋅7802 – 0⋅1238, using tables for n = 30, p = 0⋅35
- = 0⋅6564 to 4 D.P. from tables
- Example:
- A bag contains a large number of red and white discs, of which 85% are red. 20 discs are
- taken from the bag; find the probability that the number of red discs lies between 12 and 17
- inclusive.
- Solution: As there is a large number of discs in the bag, we can assume that the probability of a red
- disc remains the same for each trial, p = 0⋅85.
- Let X be the number of red discs ⇒ X ~ B(20, 0⋅85)
- We now want P(12 ≤ X ≤ 17).
- At first glance this looks simple until we realise that the tables stop at probabilities of 0⋅5.
- We need to consider the number of white discs, Y ~ B(20, 0⋅15), where 0⋅15 = 1 – 0⋅85,
- For 12 ≤ X ≤ 17
- we have X = 12, 13, 14, 15, 16 or 17
- Y = 8, 7, 6, 5, 4 or 3
- ⇒ P(12 ≤ X
-
- If X ~ B(n, p)
- then
- the expected mean is E[X] = μ = np,
- the expected variance is Var[X] = σ
- 2
- = npq = np(1 – p).
- Example: A coin is spun 100 times.
- Find the expected mean and variance of the number of heads.
- Solution: X ∼ B(100, ଵ
- ଶ
- )
- ⇒ μ = np = 100 × 1/2 = 50
- and σ
- 2
- = npq = 100 × 1/2 × 3/4 = 37⋅5
- ⇒ σ = √37 · 5 = 6⋅1237…
- μ – 2σ = 50 – 12⋅24… = 38,
- μ + 2σ = 50 + 12⋅24… = 62
- So we would expect that a probability of about 0⋅95 that the number of heads lies between 38 and 62,
- inclusive. This assumes that the probability distribution is approximately Normal.
- Example: It is believed that 35% of people like fish and chips. A survey is conducted to verify this.
- Find the minimum number of people who should be surveyed if the expected number of people
- who like fish and chips is to exceed 60.
- Solution: If X is the number of people who like fish and chips in a sample of size n,
- then X ∼ B(n, 0⋅35).
- E[x] = μ = np = 0.35n > 60
- ⇒ n > 60/0.35 = 171.4285714
- ⇒ n = 172, since the expected mean has to exceed 60.
-
- Conditions for a Poisson distribution
- Events must occur
- Singly – two cannot occur simultaneously
- Uniformly – at a constant rate
- Independently and randomly – the occurrence of one event does not influence the occurrence of
- another
-
- As before, the term probability distribution means
- the set of all possible outcomes (in this case the values of x = 0, 1, 2, …, n)
- together with their probabilities.
- Notice that in a Poisson distribution, x can take any positive or zero integral value, no matter how
- large.
- In practice, the probabilities of the ‘larger’ values will be very, very small.
- Finding probabilities for the Poisson distribution is very similar to finding probabilities for the
- Binomial
- Example: Cars pass a particular point at a rate of 5 cars per minute.
- (a) Find the probability that exactly 4 cars pass the point in a minute.
- (b) Find the probability that between at least 3 but fewer than 8 cars pass in a particular
- minute.
- (c) Find the probability that more than 8 cars pass in 2 minutes.
- (d) Find the probability that more than 3 cars pass in each of two separate minutes.
- Solution: Let X be the number of cars passing in a minute, then X ∼ PO(5)
- (a) X ∼ PO(5)
- ⇒ P(X = 4) = ହరൈ షఱ
- ସ!
- = 0⋅175467369768 = 0⋅175 to 3 S.F.
- or P(X = 4) = P(X ≤ 4) – P(X ≤ 3) = 0⋅4405 – 0⋅2650 = 0⋅1755 to 4 D.P.
- (b) P(at least 3 but fewer than 8) = P(3 ≤ X (c) We know that the Poisson distribution is uniform, so if a mean of 5 cars pass each minute,
- it means that a mean of 10 cars pass in a two minute period.
- Thus, if Y is the number of cars passing in two minutes
- Y ∼ PO(10), and we need
- P(Y > 8) = 1 – P(Y ≤ 8) = 1 – 0⋅3328 = 0⋅6672 to 4 D.P.
- (d) For probability of more than 3 in one 1 minute period, we have X ∼ PO(5)
- ⇒ P(X > 3) = 1 – P(X ≤ 2) = 1 – 0⋅1247 = 0⋅8753 from tables
- ⇒ probability of this happening in two separate minutes
- is 0⋅8753^2
- = 0⋅76615009 = 0⋅766 to 3 S.F.
-
- If X ~ PO(λ) then it can be shown that
- the expected mean is E[X] = μ = λ
- the expected variance is Var[X] = σ
- 2
- = λ.
- This means that if the set of n trials were to be repeated a large number, n, times and the number of
- occurrences recorded each time, x1, x2, x3, x4 ..... xn
- Example: In producing rolls of cloth there are on average 4 flaws in every 10 metres of cloth.
- (a) Find the mean number of flaws in a 30 metre length.
- (b) Find the probability of fewer than 3 flaws in a 6 metre length.
- (c) Find the variance of the number of flaws in a 15 metre length.
- Solution: Assuming a Poisson distribution – flaws in the cloth occur singly, independently,
- uniformly and randomly.
- (a) If the mean number of flaws in 10 metres is 4,
- then the mean number of flaws in 30 metre lengths is 3 × 4 = 12.
- (b) If there are 4 flaws on average in a 10 metre length there will be 6/10 × 4 = 2⋅4 flaws on
- average in a 6 metre length.
- If X is the number of flaws in a 6 metre length then X ∼ PO(2.4).
- P(X (c) If the mean number of flaws in 10 metre lengths is 4, then the mean number of flaws in 15
- metre lengths will be
- λ = 15/10 × 4 = 6
- Since, in a Poison distribution, the variance equals the mean
- the variance is 6.
-
- Binomial B(n, p) for small p
- If in the Binomial distribution B(n, p) p is ‘small’ and n is ‘large’, then we can approximate by a
- Poisson distribution with mean λ = np, PO(np).
- Notice that when p is ‘small’ and n is ‘large’,
- the expected variance of B(n, p) is npq ≈ np since q = 1 – p ≈ 1
- and so expected mean ≈ the variance.
- In a Poisson distribution, the expected mean = the variance, so the approximation is suitable.
- The approximation is also suitable if q is ‘small’ and n is ‘large’.
Informationen
- Beschreibung
- A-Level Maths Folienset am Statistics 2, erstellt von Sophie Edroff am 01/12/2015.
- Erstellt von
- Sophie Edroff
- Aufrufe
- 3
- Bewertung